3-10 After plotting demand for four periods, an emergency room manager has concluded that a trend-adjusted exponential smoothing model is appropriate to predict future demand. The initial estimate of trend is based on the net change of 30 for the three periods from 1 to 4, for an average of +10 units. Period Actual Period Actual 1 215 6 270 2 224 7 277 3 221 8 271 4 245 9 288 5 257 10 Use α=.5 and β=.1, and TAF of 250 for period 5. Obtain forecasts for periods 6 through 10. (Round your intermediate calculations and final answers to 2 decimal places.) t Period TAFt 6 7 8 9 10
Answer #1

EXCEL FORMULAS:

Trend Adjusted Exponentially Smoothed Forecast
Ft = TAFt-1 +α (At-1 - TAFt-1)
Tt = β(Ft – Ft-1) + (1- β) Tt-1 α = 0.5
TAFt = Ft + Tt β = 0.1
Period (t) Actual (At) Forecast (Ft) Trend (Tt) (TAFt)
1 215
2 224
3 221
4 245
5 257 =E12-D12 =(B11-B8)/3 250
6 270 =E12+$E$4*(B12-E12) =$E$5*(C13-C12)+(1-$E$5)*D12 =C13+D13
7 277 =E13+$E$4*(B13-E13) =$E$5*(C14-C13)+(1-$E$5)*D13 =C14+D14
8 271 =E14+$E$4*(B14-E14) =$E$5*(C15-C14)+(1-$E$5)*D14 =C15+D15
9 288 =E15+$E$4*(B15-E15) =$E$5*(C16-C15)+(1-$E$5)*D15 =C16+D16
10 =E16+$E$4*(B16-E16) =$E$5*(C17-C16)+(1-$E$5)*D16 =C17+D17